# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

### v2

@@ 3.1 Functional inequalities
The contraction of an outer loop of our algorithm is controlled by isoperimetry of the uniform distribution $\pi^{X}$, which is described precisely by a functional inequality. The most natural ones to consider in this setting are the Poincaré inequality (7) and log-Sobolev inequality (8). In §C, we provide a more detailed discussion of how these are related to other important notions of isoperimetry, such as the Cheeger and log-Cheeger inequalities.
@@
$\chi^{2}(\mu\mathbin{\|}\nu)\leq\frac{C_{\mathsf{PI}}(\nu)}{2}\,\mathbb{E}_{\nu}\bigl{[}\bigl{\|}\nabla\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{\|}^{2}\bigr{]}\,.$
@@
$1-\exp(-\mathcal{R}_{q}(\mu\mathbin{\|}\nu))\leq\frac{q\,C_{\mathsf{PI}}(\nu)}{4}\,\mathsf{RF}_{q}(\mu\mathbin{\|}\nu)\,,$
where $\mathsf{RF}_{q}(\mu\mathbin{\|}\nu):=q\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\lVert\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\rVert^{2}\bigr{]}\,/\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\bigr{]}$ is the Rényi Fisher information of order $q$ of $\mu$ with respect to $\nu$.
@@
We have collected below the functional inequalities used to establish the mixing of our algorithm (see §C for a detailed presentation).
@@ 3.2 Contraction and mixing
We start by analyzing how many outer iterations of $\mathsf{In\text{-}and\text{-}Out}$ are required to be $\varepsilon$-close to $\pi^{X}$, the uniform distribution over $\mathcal{K}$. The contraction of Algorithm 1 comes from analyzing Lines 2 and 3 through the perspective of heat flows (see §1.1). To exploit this view, we first revisit the previous contraction analysis in [CCSW22], which is carried out for distributions with smooth densities. Although the uniform distribution is not even continuous, we prove a technical lemma (Lemma 22) that enables us to extend previously known results to the uniform distribution. Lastly, combining the previous results with our technical lemma, we obtain clean contraction results of Algorithm 1 toward the uniform distribution $\pi^{X}$ in Theorem 23.
Part I: Contraction analysis for smooth distributions.
In this part, we review the contraction results for heat flow and its time-reversal [CCSW22], which are intimately connected with our algorithm. We also provide key technical ingredients needed for its proof, such as the computations for measures evolving under simultaneous forward/backward heat flows. We refer interested readers to §D for additional details. Only in Part I, we assume that $\nu$ denotes a probability measure with smooth density.
Forward heat flow.
We begin by introducing the “heat flow” equation (or also known as the Fokker-Planck equation), which describes the evolution of the law of $Z_{t}$ under ($\mathsf{FH}$),
$\mathbb{E}_{\mu_{h}}[f(x)]=\mathbb{E}_{\mu_{0}}[P_{h}f(x)]\,,$
where $P_{h}f(x)=\mathbb{E}_{\mathcal{N}(x,hI_{d})}[f]$.333$\{P_{h}\}_{h\geq 0}$ is often called the heat semigroup. By this we can formally identify $\mu_{h}=\mu_{0}P_{h}$, and also write $\mu_{h}$ for the law of $Z_{h}$, where $\{Z_{h}\}_{h\geq 0}$ solves ($\mathsf{FH}$).
Backward heat flow.
Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [KP21], which is the most immediately useful.
[…]
The ultimate purpose of this machinery is to affirm our earlier description of the Gibbs sampling procedure as alternating forward and backward heat flows. Indeed, notice that, if $\mu_{i}^{X}$ is the law of the iterate at some iteration $i$, then $\mu_{i}^{X}P_{h}$ is precisely $\mu_{i+1}^{Y}$ under our scheme, while $(\mu_{i}^{X}P_{h})Q_{h}^{\pi^{X},h}$ is $\mu_{i+1}^{X}$, assuming $Q_{h}^{\pi^{X},h}$ is well defined for non-smooth measures $\pi^{X}$. Thus, while Algorithm 1 is implemented via discrete steps, it can be exactly analyzed through arguments in continuous time. We shall see the benefits of this shortly.
Instead of considering the change in metrics along the evolution of $\mu P_{t}$ with respect to “fixed” $\nu$, it will be useful to consider the simultaneous evolution of $\mu P_{t},\nu P_{t}$ (and similarly $\mu Q_{t},(\nu P_{h})Q_{t}$). This type of computation was carried out for specific metrics in earlier work [VW19, CCSW22]. The following is a more generalized form of one appearing in [YFL+23, Lemma 2]. In the lemma below, we consider an arbitrary diffusion equation with corresponding Fokker-Planck equation:
$\mathrm{d}X_{t}=b_{t}(X_{t})\,\mathrm{d}t+\,\mathrm{d}B_{t}\,\quad\text{and}\quad\partial_{t}\mu_{t}=-\nabla\cdot(b_{t}\mu_{t})+\frac{1}{2}\Delta\mu_{t}\,$ (3.1)
Lemma 19 (Decay along forward/backward heat flows).
Let $(\mu_{t})_{t\geq 0},(\nu_{t})_{t\geq 0}$ denote the laws of the solutions to (3.1) starting at $\mu_{0},\nu_{0}$ respectively. Then, for any differentiable function $g$,
@@
For brevity, we drop the variable $x$ of functions involved, and proceed as follows:
$\partial_{t}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})$ =$=\int\Bigl{\{}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\partial_{t}\nu_{t}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\bigl{(}\frac{\mu_{t}}{\nu_{t}}\bigr{)}^{\prime}\nu_{t}\Bigr{\}}\,\mathrm{d}x$
=$=\int\Bigl{\{}\partial_{t}\nu_{t}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\partial_{t}\mu_{t}\Bigr{\}}\,\mathrm{d}x$
=$\underset{(i)}{=}\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})+\frac{1}{2}\Delta\nu_{t}\bigr{]}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}\,\mathrm{d}x$
where in $(i)$ we substitute the F-P equation from (3.1). Integrating by parts (i.e., $\int f\operatorname{div}(\mathbf{G})=-\int\langle\nabla f,\mathbf{G}\rangle$ for a real-valued function $f$ and vector-valued function $\mathbf{G}$), we have that
$\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})\bigr{]}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\,\mathrm{d}x=\int\Bigl{\langle}b_{t}\nu_{t},\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x\,.$ (3.2)
On the other hand, we have that
$-\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})\bigr{]}\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\,\mathrm{d}x=-\int\Bigl{\langle}b_{t}\nu_{t},\frac{\mu_{t}}{\nu_{t}}\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x\,.$
The second term cancels with the RHS of (3.2). We have a similar cancellation for the $\frac{1}{2}\Delta\nu_{t}$ term:
and
$-\int\frac{1}{2}\Delta\nu_{t}\,\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\,\mathrm{d}x=\int\frac{1}{2}\Bigl{\langle}\nabla\nu_{t},\frac{\mu_{t}}{\nu_{t}}\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x\,.$
Combining these, we are left with
$\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})+$ $\frac{1}{2}\Delta\nu_{t}\bigr{]}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}\,\mathrm{d}x$
$=-\int\Bigl{\langle}b_{t}\nu_{t}-\frac{1}{2}\nabla\nu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x$
Finally, we note that
which completes the proof. ∎
Before proceeding, we need a standard lemma on functional inequalities under ($\mathsf{FH}$).
The following inequality holds for any $\pi$ with finite log-Sobolev and Poincaré constants,
$C_{\mathsf{PI}}(\pi P_{t})\leq C_{\mathsf{PI}}(\pi)+t\,,\qquad\text{and}\qquad C_{\mathsf{LSI}}(\pi P_{t})\leq C_{\mathsf{LSI}}(\pi)+t\,.$
Combining the previous two lemmas, we can establish contraction between $\mu P_{h}Q_{h}$ and $\nu$ after one forward/backward iteration.
Theorem 21 ([CCSW22, Theorem 3 and 4]).
Assume $\nu$, a measure with smooth density, satisfies (8) with constant $C_{\mathsf{LSI}}$. For any $q\geq 1$ and initial distribution $\mu$ with a smooth density, denoting again $Q_{h}:=Q_{h}^{\nu,h}$,
$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{LSI}})^{2/q}}\,.$
If $\nu$ satisfies (7) with constant $C_{\mathsf{PI}}$, then it follows that
$\chi^{2}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\frac{\chi^{2}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{PI}})^{2}}\,.$
Moreover, for all $q\geq 2$,
$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\begin{cases}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)-\frac{2\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{PI}})^{2/q}}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
Proof.
Since the SDE in (3.1) captures the forward heat flow ($\mathsf{FH}$), we set $\mu_{0}$ and $\nu_{0}$ in Lemma 19 to $\mu$ and $\nu$, respectively, obtaining contraction along the forward heat flow as follows: Substituting the $q$-Rényi into Lemma 19, we have, from the definition of the Rényi divergence as $\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{1}{q-1}\log\mathopen{}\mathclose{{}\left(D_{f}(\mu\mathbin{\|}\nu)+1}\right),$ with $f(x)=x^{q}-1$ and $g(x)=\frac{1}{q-1}\log(x+1)$,
$\partial_{t}\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})$ $=-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{\langle}\nabla\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-1},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\Bigr{\rangle}\Bigr{]}}{(q-1)\,\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}$
=$\underset{(i)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-2}\langle\nabla\frac{\mu P_{t}}{\nu P_{t}},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\rangle\Bigr{]}}{\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}$
=$\underset{(ii)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}\lVert\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\rVert^{2}}\right]}{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}}\right]}=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})\,,$
where in $(i)$, we use again that $\nabla\bigl{[}f^{\prime}\bigl{(}\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\bigr{]}=\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\cdot\frac{\mu_{t}}{\nu_{t}}+f^{\prime}\bigl{(}\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}$, and $(ii)$ uses that $\nabla\frac{\mu P_{t}}{\nu P_{t}}=\frac{\mu P_{t}}{\nu P_{t}}\nabla\log\frac{\mu P_{t}}{\nu P_{t}}$, and the last equality recalls the definition of the Rényi Fisher information. This yields
$\partial_{t}\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})=-\frac{1}{2}\mathsf{RF}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})\underset{(i)}{\leq}-\frac{1}{q}\frac{\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{LSI}}(\nu P_{t})}\underset{(ii)}{\leq}-\frac{1}{q}\frac{\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{LSI}}+t}\,,$
@@
where we used Lemma 17 in $(i)$ and Lemma 20 in $(ii)$. Applying Grönwall’s inequality (Lemma 10),
$\mathcal{R}_{q}(\mu P_{h}\mathbin{\|}\nu P_{h})\leq\exp\Bigl{(}-\frac{1}{q}\int_{0}^{h}\frac{1}{C_{\mathsf{LSI}}+t}\,\mathrm{d}t\Bigr{)}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{LSI}})^{1/q}}\,.$
Since the SDE (3.1) also captures the backward equation ($\mathsf{BH}$), we set $\mu_{0}$ and $\nu_{0}$ in Lemma 19 to $\mu P_{h}$ and $\tilde{\nu}:=\nu P_{h}$ respectively, obtaining contraction along the backward heat flow:
$\partial_{t}\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})$ $=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})$
$\leq-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}(\tilde{\nu}Q_{t})}\underset{(i)}{\leq}-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}+{h-t}}\,,$
where $(i)$ follows from that $\tilde{\nu}Q_{t}=\nu P_{h}Q_{t}=\nu P_{h-t}$ and $C_{\mathsf{LSI}}(\tilde{\nu}Q_{t})\leq C_{\mathsf{LSI}}+h-t$ due to Lemma 20. Applying Lemma 10 again yields
$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\frac{\mathcal{R}_{q}(\mu P_{h}\mathbin{\|}\tilde{\nu})}{(1+h/C_{\mathsf{LSI}})^{1/q}}\,.$
Composing these two inequalities leads to the decay rate claimed in the theorem.
@@
$\partial_{t}\chi^{2}(\mu P_{t}\mathbin{\|}\nu P_{t})=-\frac{1}{2}\,\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{\|}\nabla\frac{\mu P_{t}}{\nu P_{t}}\Bigr{\|}^{2}\Bigr{]}\underset{(i)}{\leq}-\frac{\chi^{2}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{PI}}(\nu P_{t})}\leq-\frac{\chi^{2}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{PI}}+t}\,,$
@@
$\chi^{2}(\mu P_{h}\mathbin{\|}\nu P_{h})\leq\exp\Bigl{(}-\int_{0}^{h}\frac{1}{C_{\mathsf{PI}}+t}\,\mathrm{d}t\Bigr{)}\,\chi^{2}(\mu\mathbin{\|}\nu)\leq\frac{\chi^{2}(\mu\mathbin{\|}\nu)}{1+h/C_{\mathsf{PI}}}\,.$
The decay along the backward heat flow in $\chi^{2}$ is entirely analogous to the Rényi case. Then we combine two contraction results from the forward and backward flows, completing the proof.
@@
The result in the $\mathcal{R}_{q}$ under (7) can be shown in a similar manner. The only difference is that in forward and backward computations, one should use the functional inequality in Lemma 16 and the following standard inequalities:

### v3

@@ 3.1 Functional inequalities
The contraction of an outer loop of our algorithm is controlled by isoperimetry of the uniform distribution $\pi^{X}$, which is described precisely by a functional inequality. The most natural ones to consider in this setting are the Poincaré inequality (7) and log-Sobolev inequality (8). In §B, we provide a more detailed discussion of how these are related to other important notions of isoperimetry, such as the Cheeger and log-Cheeger inequalities.
@@
$\chi^{2}(\mu\mathbin{\|}\nu)\leq\frac{C_{\mathsf{PI}}(\nu)}{2}\,\mathbb{E}_{\nu}\bigl[\big\lvert\nabla\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\big\rvert^{2}\bigr]\,.$
@@
$1-\exp\bigl(-\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\bigr)\leq\frac{q\,C_{\mathsf{PI}}(\nu)}{4}\,\mathsf{RF}_{q}(\mu\mathbin{\|}\nu)\,$,
where $\mathsf{RF}_{q}(\mu\mathbin{\|}\nu):=q\,\mathbb{E}_{\nu}[(\frac{\mathrm{d}\mu}{\mathrm{d}\nu})^{q}\,\lvert\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\rvert^{2}]\,/\,\mathbb{E}_{\nu}[(\frac{\mathrm{d}\mu}{\mathrm{d}\nu})^{q}]$ is the Rényi Fisher information of order $q$ of $\mu$ with respect to $\nu$.
@@
We have collected below the functional inequalities used to establish the mixing of our algorithm (see §B for a detailed presentation).
@@ 3.2 Contraction and mixing
We start by analyzing how many outer iterations of $\mathsf{In\text{-}and\text{-}Out}$ are required to be $\varepsilon$-close to $\pi^{X}$, the uniform distribution over $\mathcal{K}$. The contraction of Algorithm 1 comes from analyzing Lines 2 and 3 through the perspective of heat flows (see §1.1). Unlike prior works on uniform sampling, we do not bound the $s$-conductance of this chain. Instead, we use the “calculus” of the space of probability measures; i.e., for some probability divergence $D$, we compute $\partial_{t}D(\mu*\gamma_{t}\mathbin{\|}\pi^{X}*\gamma_{t})$ in terms of $t$ and isoperimetric constants of $\pi^{X}$.
The classical data processing inequality (DPI; see Lemma 9) shows that $f$-divergence between two distributions cannot increase when both are convolved with a Gaussian. To prove contraction, we will use a strong data-processing inequality (SDPI) for $f$-divergences, which gives a quantitative bound on the contraction. Such an inequality was known for some cases including the KL-divergence [Sta59], and was proven for $f(x)=x^{q}-1$ by Klartag and Ordentlich [KO25] who also established the SDPI more generally for this type of Gaussian convolution [AG76, PW16, CPW18], showing that finite fourth moment (i.e., $\mathbb{E}_{\pi^{X}}[\lvert\cdot\rvert^{4}]<\infty$) is a sufficient condition. In §C, we show how the identity is derived under suitable regularity assumptions (see Lemma 28).
Proposition 1 (General de Bruijn identity).
Let $\mu$ and $\nu$ be probability measures on $\mathbb{R}^{d}$ such that $\mathbb{E}_{\nu}[\lvert\cdot\rvert^{4}]<\infty$ and $\chi^{q}(\mu\mathbin{\|}\nu)<\infty$. Then, for any $t>0$ and $q>1$,
$\partial_{t}\chi^{q}(\mu_{t}\mathbin{\|}\nu_{t})=-\frac{q(q-1)}{2}\,\mathbb{E}_{\nu_{t}}\bigl[\bigl(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\bigr)^{q}\bigl|\nabla\log\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\bigr|^{2}\bigr]=-\frac{q-1}{2}\,\Bigl\|\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\Bigr\|^{q}_{L^{q}(\nu_{t})}\mathsf{RF}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\,.$
Using the general de Bruijn identity with $\mathcal{R}_{q}=\frac{1}{q-1}\log(1+\chi^{q})$, the chain rule results in
$\partial_{t}\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})=\frac{1}{q-1}\,\frac{\partial_{t}\chi^{q}(\mu_{t}\mathbin{\|}\nu_{t})}{1+\chi^{q}(\mu_{t}\mathbin{\|}\nu_{t})}=\frac{1}{q-1}\,\frac{\partial_{t}\chi^{q}(\mu_{t}\mathbin{\|}\nu_{t})}{\lVert\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\rVert^{q}_{L^{q}(\nu_{t})}}=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\,.$ (3.1)
We also obtain $\partial_{t}\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})=-\mathbb{E}_{\nu_{t}}[\lvert\nabla\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\rvert^{2}]$.
Before proceeding, we need a property of the stability of functional inequalities under the heat flow.
The following inequality holds for any $t>0$ and $\pi$ with finite log-Sobolev and Poincaré constants,
$C_{\mathsf{PI}}(\pi_{t})\leq C_{\mathsf{PI}}(\pi)+t\,,\qquad\text{and}\qquad C_{\mathsf{LSI}}(\pi_{t})\leq C_{\mathsf{LSI}}(\pi)+t\,.$
We can now deduce contraction under the heat flow.
Lemma 19 (Contraction under functional inequalities).
Let $\mu,\nu$ be probability measures on $\mathbb{R}^{d}$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\nu)<\infty$. Then, for any $t>0$ and $q>1$,
$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+t/C_{\mathsf{LSI}}(\nu))^{1/q}}\,.$
Also, $\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})\leq\frac{\chi^{2}(\mu\mathbin{\|}\nu)}{(1+t/C_{\mathsf{PI}}(\nu))}$. Furthermore, for any $q\geq 2$,
$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\begin{cases}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)-\frac{\log(1+t/C_{\mathsf{PI}}(\nu))}{q}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+t/C_{\mathsf{PI}}(\nu))^{1/q}}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)<1\,.\end{cases}$
@@
It is well-known that (8) implies (7), and that (7) ensures finite moments of any order due to exponential integrability (see [BGL14, Proposition 4.4.2]). Also, both $\mu_{t}$ and $\nu_{t}$ are smooth due to the Gaussian convolution.
Under $C_{\mathsf{LSI}}(\nu)<\infty$, (3.1) implies that
$\partial_{t}\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\underset{(i)}{\leq}-\frac{\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})}{q\,C_{\mathsf{LSI}}(\nu_{t})}\underset{(ii)}{\leq}-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})}{C_{\mathsf{LSI}}(\nu)+t}\,$,
@@
where we used Lemma 16 in $(i)$ and Lemma 18 in $(ii)$. Applying Grönwall’s inequality (Lemma 10), we deduce that
$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\exp\Bigl(-\frac{1}{q}\int_{0}^{t}\frac{1}{C_{\mathsf{LSI}}(\nu)+s}\,\mathrm{d}s\Bigr)\,\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+t/C_{\mathsf{LSI}}(\nu))^{1/q}}\,.$
@@
$\partial_{t}\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})=-\frac{1}{2}\,\mathbb{E}_{\nu_{t}}\bigl[\big\lvert\nabla\frac{\mu_{t}}{\nu_{t}}\big\rvert^{2}\bigr]\underset{(i)}{\leq}-\frac{\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})}{C_{\mathsf{PI}}(\nu_{t})}\leq-\frac{\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})}{C_{\mathsf{PI}}(\nu)+t}\,$,
@@
$\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})\leq\exp\Bigl(-\int_{0}^{t}\frac{1}{C_{\mathsf{PI}}(\nu)+s}\,\mathrm{d}s\Bigr)\,\chi^{2}(\mu\mathbin{\|}\nu)\leq\frac{\chi^{2}(\mu\mathbin{\|}\nu)}{1+t/C_{\mathsf{PI}}(\nu)}\,.$
@@
The result in the $\mathcal{R}_{q}$ under (7) can be shown in a similar manner. The only difference is that in forward computation, one should use the functional inequality in Lemma 15 and the following standard inequalities: